OpenStudy (anonymous):

lim x->0^+ (sinx)/(x+sqrootx) Answer: 0 My question, why 0? Shouldn't it be infinity?

OpenStudy (anonymous):

if x = 0 this equals 0 / 0 so use l'hopitals rule differentiate top and bottom

OpenStudy (anonymous):

i'll take a look at it - but limits are not my strong point

OpenStudy (anonymous):

So I used l'hopitals rule and I keep getting 0/0

OpenStudy (anonymous):

heres the man - zarkon

OpenStudy (anonymous):

:D

OpenStudy (zarkon):

look at it like this... \[\frac{\sin(x)}{x+\sqrt{x}}=\frac{1}{\frac{x+\sqrt{x}}{\sin{x}}}=\frac{1}{\frac{x}{\sin(x)}+\frac{\sqrt{x}}{\sin{x}}}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0^+}\frac{\sin(x)}{x+\sqrt{x}} \cdot \frac{x- \sqrt{x}}{x-\sqrt{x}}=\lim_{x \rightarrow 0^+}\frac{\sin(x)(x-\sqrt{x})}{x^2-x}\] \[\lim_{x \rightarrow 0^+}\frac{\sin(x)(x-\sqrt{x})}{x(x-1)}=\lim_{x \rightarrow 0^+}\frac{\sin(x)}{x} \cdot \lim_{x \rightarrow 0^+} \frac{x-\sqrt{x}}{x-1}\]

myininaya (myininaya):

choose my way :)

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

haha thanks :D, so at the end what should i replace x as? A number close to 0? The answer they give me is 0 :S

myininaya (myininaya):

\[\lim_{x \rightarrow 0^+}\frac{\sin(x)}{x}=1\] \[\lim_{x \rightarrow 0^+}\frac{x-\sqrt{x}}{x-1}=\frac{0-\sqrt{0}}{0-1}=0\] the answer is 0*1 which can be simplify to ..... you fill in this blank lol

OpenStudy (anonymous):

Ohhhhhhhh, I get it now.. I guess I should be more careful next time. Thanks a lot , I appreciate it.